# Zero-Waiting Load Balancing with Heterogeneous Servers in Heavy Traffic

### v1

@@ 3.3 Generator Approximation via Stein’s Method
To analyze the steady-state delay performance of the load-balancing system, we adopt the generator approximation method via Stein’s method. This technique allows us to couple the original high-dimensional stochastic system with a simpler, low-dimensional reference system, whose behavior is more tractable. The generator discrepancy between these two systems determines the approximation error of the target metric, which is a test function of queue lengths. Following [21], we choose the test function to be the clipped $r$-th moment of average queue length at steady-state. Recall that the most of the fastest servers are busy whenever there are waiting jobs in the system in Lemmas 3 and 4. This motivates us to consider a simplified single-queue system where the service rate is slightly above the arrival rate. Specifically, the fluid system evolves according to the ODE
@@
Our interesting metric is the moment bound of total queue length, and we consider the $r$-th moment of the test function
@@
Let $G$ denote the generator of the original queueing process under load balancing algorithms, we have the Stein identity:
@@
The full derivation is included in Appendix E for completeness. If we ignore the term in (11), we already have the iteration between the $r$th-order moment and the (r1)(r-1)th-order moment of queue length. Therefore, the key is to quantify the term in (11) and $(\lambda+\frac{1}{N^{\alpha}}-\frac{1}{N}\sum_{n=1}^{N}\mu_{n}\mathbb{I}(Q_{n}\geq 1))$ in particular. Intuitively, when the average queue length is greater than a threshold $\bar{Q}>\eta+\frac{1}{N}$} and almost all servers are busy under JFSQ, the average service rate $\frac{1}{N}\sum_{n=1}^{N}\mu_{n}\mathbb{I}(Q_{n}\geq 1)$ should be “close” to $\lambda+\frac{1}{N^{\alpha}}=1.$ Specifically, if $\frac{1}{N}\sum_{n=1}^{N}\mu_{n}\mathbb{I}(Q_{n}\geq 1)\geq 1-\frac{c}{N^{\alpha}}$ holds for a constant smaller than $c<1$. The term in (11) can be absorbed by the metric $\mathbb{E}\left[h^{r}\left(\frac{1}{N}\sum_{n=1}^{N}Q_{n}\right)\right]$ and we again have a clean iteration for the $r$th-order moment and the (r1)(r-1)th-order moment of queue length. Next, we justify the intuition for sub and super Haffin-Whitt regimes, where we use distinct thresholds $\eta$ for both regimes. We first prove the high-order moment bounds in Theorem 2 and then asymptotic zero-waiting results in Theorem 1.
@@ 3.4 Proving Theorem 2
Now define event $\tilde{\cal E}_{sub}:=\mathcal{E}_{1}\cap\mathcal{E}_{sub}.$ For the system states in $\tilde{\cal E}_{sub},$ we have
@@
$\sum_{n=1}^{N}(Q_{n}-1)\mathbb{I}({Q_{n}\geq 2})=\sum_{n=1}^{N}Q_{n}-\sum_{n=1}^{N}\mathbb{I}({Q_{n}\geq 1})>N_{2}+(1-\frac{1}{4\nu})N^{1-\alpha}-\sum_{n=1}^{N}\mathbb{I}({Q_{n}\geq 1}).$
Therefore, when the event $\tilde{\cal E}_{sub}$ happens, we have
@@
where the first inequality holds due to the definition of $\sum_{n=1}^{N_{2}}\mu_{n}=\lambda N$ and the inequality (15); the second inequality holds according to Lemmas 1 and 2; the last inequality holds because of the assumption $N^{1-\alpha}\geq\frac{16\nu}{\epsilon}(\mu_{N_{2}}+\mu_{1}\left(2\log^{2}N+1\right)+3\mu_{N_{1}}\sqrt{N}\log N).$ Therefore, when the event $\tilde{\mathcal{E}}_{sub}$ happens, we have
@@
$\leq$ $2b^{r}N^{\alpha}\Pr\left(\mathcal{E}_{sub}^{c}\right)+\frac{2^{r+2}}{N^{r-\alpha}}+\frac{r}{N^{1-\alpha}}\mathbb{E}\left[{h^{r-1}_{k}}\left(\frac{1}{N}\sum_{n=1}^{N}Q_{n}+{\frac{1}{N}}\right)\right]$
$\leq$ $\frac{2^{r+2}+2}{N^{r-\alpha}}+\frac{r}{N^{1-\alpha}}\mathbb{E}\left[{h^{r-1}_{k}}\left(\frac{1}{N}\sum_{n=1}^{N}Q_{n}+{\frac{1}{N}}\right)\right],$
where the last inequality holds by $\mathbb{P}(\mathcal{E}_{sub}^{c})\leq\frac{1}{N^{2r}}$ and $\frac{b^{r}}{N^{r-\alpha}}\leq 2.$ Recall $k_{sub}=\frac{16\nu}{\epsilon}=\frac{16(b-1)\mu_{N_{2}}}{\epsilon^{2}}$, we have:
@@
where the last inequality holds because of the definition of $\sum_{n=1}^{N}\mu_{n}=N$ and the assumption of $\mu_{1}\geq\cdots\geq\mu_{n}$. Therefore, when $\tilde{\mathcal{E}}_{super}$ happens we have
$\lambda+\frac{1}{N^{\alpha}}-\frac{1}{N}\left(\sum_{n=1}^{N}\mu_{n}\mathbb{I}(Q_{n}\geq 1)\right)\leq\frac{1}{2N^{\alpha}}.$ (24)
@@
where the last inequality holds by $\mathbb{P}(\tilde{\cal E}^{c})\leq\frac{1}{N^{2r}}$ according to Lemma 4 and $\frac{b^{r}}{N^{r-\alpha}}\leq 2.$ Now we have established the iterative relationship of high-order moments. The remaining part follows the same proof by expanding (25), as in Sub-Haffin-Whitt regime in (17)–(21).
@@ 3.5 Proving Theorem 1
$p_{\mathcal{B}}\leq$ $\mathbb{P}\left(\sum_{n=1}^{N}Q_{n}\geq Nb\right)\leq P\left(h^{r}(\bar{Q})\geq(b-\eta)^{r}\right)$
@@
$\mathbb{E}\left[\sum_{n=1}^{N}Q_{n}\right]\leq N_{2}+(34k_{sub}+\log N)N^{\alpha}.$
@@
$\leq$ $N_{2}+(34k_{sub}+\log N)N^{\alpha}-\mathbb{E}\left[\sum_{n=1}^{N_{2}}\mathbb{I}(Q_{n}\geq 1)\right]$
$=$ $(34k_{sub}+\log N)N^{\alpha}+\mathbb{E}\left[\sum_{n=1}^{N_{2}}\mathbb{I}(Q_{n}=0)\right]$
$\leq$ $(34k_{sub}+\log N)N^{\alpha}+6\sqrt{N}\log N$
@@
$\mathbb{E}[W]\leq$ $\frac{(34k_{\rm sub}+\log N)N^{\alpha}+6\sqrt{N}\log N}{\lambda N(1-p_{\mathcal{B}})}.$
@@
Waiting probability: Recall $\overline{\mathcal{W}}$ to be the waiting event that a job is not blocked (i.e., a job is routed to a busy server and waits in a buffer) and $p_{\overline{\mathcal{W}}}$ to be its steady-state probability. Let $T_{Q}$ be the queueing time for the jobs waiting in the buffer and $S$ be the service time. According to Little’s law, we have
@@
$p_{\overline{\mathcal{W}}}\leq$ $(1-p_{\mathcal{B}})\frac{(34k_{sub}+\log N)N^{\alpha}+6\sqrt{N}\log N}{(1-\frac{1}{N})(N_{2}-5\sqrt{N}\log N)}$
$\leq$ $\frac{(34k_{sub}+\log N)N^{\alpha}+6\sqrt{N}\log N}{N_{2}-5\sqrt{N}\log N}$
$\leq$ $\frac{(34k_{sub}+\log N)N^{\alpha}+6\sqrt{N}\log N}{N-\frac{N^{1-\alpha}}{\epsilon}-5\sqrt{N}\log N}$
$\leq$ $\frac{4(34k_{sub}+\log N)N^{\alpha}+24\sqrt{N}\log N}{N}$
where the second inequality holds because $p_{\mathcal{B}}\leq 1/N$; the third inequality holds because the lower bound of $N_{2}$ is obtained when the service rate $\mu_{n}=\epsilon,\forall n>N_{2}$, i.e., $(N-N_{2})\epsilon=N^{1-\alpha}$; the last inequality holds for a large $N$ such that $\epsilon\geq 2/N^{\alpha}$ and $\sqrt{N}\geq 20\log N$. Finally, a job that is not routed to an idle server is either blocked or wait in a buffer
$p_{\mathcal{W}}=$ $p_{\mathcal{B}}+p_{\overline{\mathcal{W}}}\leq\frac{4(34k_{sub}+\log N)N^{\alpha}+24\sqrt{N}\log N+1}{N}$
@@
$p_{\mathcal{B}}\leq$ $\mathbb{P}\left(\sum_{n=1}^{N}Q_{n}\geq Nb\right)\leq 32\left(\frac{2r}{N^{(1-\alpha)}}\right)^{r}.$
@@ Appendix B Most of the Fastest Servers are Busy
$\Pr\left(V(q)\geq 1+8\log^{2}N\right)$ $\leq\left(\frac{N}{N+\frac{\delta N}{2}}\right)^{4\log^{2}N}$
@@
$\Pr\left(V(q)\geq 5\sqrt{N}\log N\right)\leq$ $\left(\frac{N}{N+\mu_{N_{2}}{\sqrt{N}}\log N}\right)^{2\sqrt{N}\log N}$
@@ Appendix C State Space Collapse in Sub-Halfin-Whitt Regime
$=$ $\sum_{n=1}^{N}\lambda NA_{n}(q)(V(q+e_{n})-V(s))+\mu_{n}\mathbb{I}(q_{n}\geq 1)(V(q-e_{n})-V(q)).$
@@
    $=$ $\min\left\{\sum_{n=1}^{N}(q_{n}-1)\mathbb{I}(q_{n}\geq 2)-kN^{\alpha}\log N,~N_{2}+(1-\frac{1}{4\nu})N^{1-\alpha}-\sum_{n=1}^{N}\mathbb{I}(q_{n}\geq 1)-1\right\}-V(q)$
@@
    $\triangledown V(s)\leq$ $-\lambda N+\sum_{m=1}^{N}\mu_{m}\mathbb{I}(q_{m}=1)$
@@
Now invoking Lemma 5 with $B=c_{0}N^{1-\alpha},$ $\gamma=\frac{\epsilon}{2(b-1)}N^{1-\alpha},$ $\nu_{\max}=1,$ $q_{\max}\leq N,$ $\delta=N$ and $j=\frac{1-c_{0}}{4}N^{1-\alpha},$ we obtain that for sufficiently large NN
$\Pr\left(V(q)\geq\frac{1+c_{0}}{2}N^{1-\alpha}\right)$ $\leq\left(\frac{N}{N+\frac{\epsilon}{4(b-1)}N^{1-\alpha}}\right)^{\frac{1-c_{0}}{4}N^{1-\alpha}}+\left(\frac{N}{\frac{\epsilon}{4(b-1)}N^{1-\alpha}}+1\right)\exp\left(-\epsilon\log^{2}N\right)$

### v2

@@ 3.3 Generator Approximation via Stein’s Method
To analyze the steady-state delay performance of the load-balancing system, we adopt the generator approximation method via Stein’s method. This technique allows us to couple the original high-dimensional stochastic system with a simpler, low-dimensional reference system, whose behavior is more tractable. The generator discrepancy between these two systems determines the approximation error of the target metric, which is a test function of queue lengths. Following [21], we choose the test function to be the clipped $r$-th moment of average queue length at steady-state. Recall that the most of the fastest servers are busy whenever there are waiting jobs in the system in Lemmas 3 and 4. This motivates us to consider the fluid model of a simplified single-queue system where the service rate is slightly above the arrival rate. Specifically, the fluid system evolves according to the ODE
@@
Our metric of interesting is the moment bound of total queue length, and we consider the $r$-th moment of the test function
@@
Let $G$ denote the generator of the original queueing process under load balancing algorithms. $g(\bar{Q})$ is bounded according to the (7) and the finite buffer assumptioni.e., $\bar{Q}\leq b$, we have the Stein identity:
@@
The full derivation is included in Appendix E for completeness. If we ignore the term in (11), we already have the iteration between the $r$th-order moment and the (r1)(r-1)th-order moment of queue length. Therefore, the key is to quantify the term in (11) and $(\lambda+\frac{1}{N^{\alpha}}-\frac{1}{N}\sum_{n=1}^{N}\mu_{n}\mathbb{I}(Q_{n}\geq 1))$ in particular. Intuitively, when the average queue length is greater than a threshold $\bar{Q}>\eta+\frac{1}{N}$} and almost all servers are busy under JFSQ, the average service rate $\frac{1}{N}\sum_{n=1}^{N}\mu_{n}\mathbb{I}(Q_{n}\geq 1)$ should be “close” to $\lambda+\frac{1}{N^{\alpha}}=1.$ Specifically, if $\frac{1}{N}\sum_{n=1}^{N}\mu_{n}\mathbb{I}(Q_{n}\geq 1)\geq 1-\frac{c}{N^{\alpha}}$ holds for a constant smaller than $c<1$, we have the key term in (11)
$\mathbb{E}\left[N^{\alpha}h^{r}\left(\bar{Q}\right)\left(\lambda+\frac{1}{N^{\alpha}}-\frac{1}{N}\sum_{n=1}^{N}\mu_{n}\mathbb{I}(Q_{n}\geq 1)\right)\mathbb{I}({\bar{Q}>\eta+\frac{1}{N}})\right]\leq c\cdot\mathbb{E}\left[h^{r}\left(\bar{Q}\right)\right].$
This term can be absorbed into $\mathbb{E}\left[h^{r}\left(\bar{Q}\right)\right]$ by moving it to the left-hand-side, resulting in
$(1-c)\mathbb{E}\left[h^{r}\left(\bar{Q}\right)\right]\leq\frac{2^{r+2}}{N^{r-\alpha}}+\frac{r}{N^{1-\alpha}}\mathbb{E}\left[h^{r-1}\left(\bar{Q}+\frac{1}{N}\right)\right].$
We again have a clean iteration for the $r$th-order moment and the (r1)(r-1)th-order moment of queue length. Next, we justify the intuition for sub and super Haffin-Whitt regimes, where we use distinct thresholds $\eta$ for both regimes. We first prove the high-order moment bounds in Theorem 2 and then asymptotic zero-waiting results in Theorem 1.
@@ 3.4 Proving Theorem 2
For the system states $Q\in\mathcal{E}_{1}\cap\mathcal{E}_{sub},$ we have
@@
$\sum_{n=1}^{N}(Q_{n}-1)\mathbb{I}({Q_{n}\geq 2})=$ $\sum_{n=1}^{N}Q_{n}-\sum_{n=1}^{N}\mathbb{I}({Q_{n}\geq 1})$
$>$ $N_{2}+(1-\frac{1}{4\nu})N^{1-\alpha}-\sum_{n=1}^{N}\mathbb{I}({Q_{n}\geq 1}).$
Therefore, when the system states $Q\in\mathcal{E}_{1}\cap\mathcal{E}_{sub}$, we have
@@
where the first inequality holds due to the definition of $\sum_{n=1}^{N_{2}}\mu_{n}=\lambda N$ and the inequality (15); the second inequality holds according to Lemmas 1 and 2; the last inequality holds because Assumption 1 implies $N^{1-\alpha}\geq\frac{16\nu}{\epsilon}(\mu_{N_{2}}+\mu_{1}\left(2\log^{2}N+1\right)+3\mu_{N_{1}}\sqrt{N}\log N)$. Therefore, when the event $\tilde{\mathcal{E}}_{sub}$ happens, we have
@@
$\leq$ $2b^{r}N^{\alpha}\mathbb{P}\left(\mathcal{E}_{sub}^{c}\right)+\frac{2^{r+2}}{N^{r-\alpha}}+\frac{r}{N^{1-\alpha}}\mathbb{E}\left[{h^{r-1}}\left(\frac{1}{N}\sum_{n=1}^{N}Q_{n}+{\frac{1}{N}}\right)\right]$
$\leq$ $\frac{2^{r+2}+2}{N^{r-\alpha}}+\frac{r}{N^{1-\alpha}}\mathbb{E}\left[{h^{r-1}}\left(\frac{1}{N}\sum_{n=1}^{N}Q_{n}+{\frac{1}{N}}\right)\right],$
where the last inequality holds by $\mathbb{P}(\mathcal{E}_{sub}^{c})\leq\frac{1}{N^{2r}}$ in Lemma 3 and $\frac{b^{r}}{N^{r-\alpha}}\leq 2$ in Assumption 1. Recall $k_{sub}=\frac{16\nu}{\epsilon}=\frac{16(b-1)\mu_{N_{2}}}{\epsilon^{2}}$, we have:
@@
where the last inequality holds because of the definition of $\sum_{n=1}^{N}\mu_{n}=N$ and the assumption of $\mu_{1}\geq\cdots\geq\mu_{n}$. Therefore, recalling that $\lambda+\frac{1}{N^{\alpha}}=1$, on the event $\tilde{\mathcal{E}}_{\mathrm{super}}$ we have
$\lambda+\frac{1}{N^{\alpha}}-\frac{1}{N}\left(\sum_{n=1}^{N}\mu_{n}\mathbb{I}(Q_{n}\geq 1)\right)\leq\lambda+\frac{1}{N^{\alpha}}-\frac{1}{N}(N-\frac{1}{2}N^{1-\alpha})=\frac{1}{2N^{\alpha}}.$ (24)
@@
where the last inequality holds by $\mathbb{P}(\tilde{\cal E}^{c})\leq\frac{1}{N^{2r}}$ in Lemma 4 and $\frac{b^{r}}{N^{r-\alpha}}\leq 2$ in Assumption 2. Now we have established the iterative relationship of high-order moments. The remaining part follows the same proof by expanding (25), as in Sub-Haffin-Whitt regime in (17)–(21).
@@ 3.5 Proving Theorem 1
$p_{\mathcal{B}}\leq$ $\mathbb{P}\left(\sum_{n=1}^{N}Q_{n}=Nb\right)\leq P\left(h^{r}(\bar{Q})\geq(b-\eta)^{r}\right)$
@@
$\mathbb{E}\left[\sum_{n=1}^{N}Q_{n}\right]\leq N_{2}+N^{1-\alpha}+(34k_{sub}+\log N)N^{\alpha}.$
@@
$\leq$ $N_{2}+N^{1-\alpha}+(34k_{sub}+\log N)N^{\alpha}-\mathbb{E}\left[\sum_{n=1}^{N_{2}}\mathbb{I}(Q_{n}\geq 1)\right]$
$=$ $N^{1-\alpha}+(34k_{sub}+\log N)N^{\alpha}+\mathbb{E}\left[\sum_{n=1}^{N_{2}}\mathbb{I}(Q_{n}=0)\right]$
$\leq$ $N^{1-\alpha}+(34k_{sub}+\log N)N^{\alpha}+6\sqrt{N}\log N$
@@
$\mathbb{E}[W]\leq$ $\frac{N^{1-\alpha}+(34k_{\rm sub}+\log N)N^{\alpha}+6\sqrt{N}\log N}{\lambda N(1-p_{\mathcal{B}})}.$
@@
Waiting probability: Recall $\overline{\mathcal{W}}$ to be the waiting event that a job is not blocked (i.e., a job is routed to a busy server and waits in a buffer) and $p_{\overline{\mathcal{W}}}$ to be its steady-state probability. Let $T_{Q}$ be the queueing time for the jobs waiting in the buffer (i.e., the waiting time of jobs not receiving service immediately) and $S$ be the service time of jobs in service. According to Little’s law (Little’s Law for Time in Queue of Corollary 6.4 in [15]), we have
@@
$p_{\overline{\mathcal{W}}}\leq$ $(1-p_{\mathcal{B}})\frac{N^{1-\alpha}+(34k_{sub}+\log N)N^{\alpha}+6\sqrt{N}\log N}{(1-\frac{1}{N})(N_{2}-5\sqrt{N}\log N)}$
$\leq$ $\frac{N^{1-\alpha}+(34k_{sub}+\log N)N^{\alpha}+6\sqrt{N}\log N}{N_{2}-5\sqrt{N}\log N}$
$\leq$ $\frac{N^{1-\alpha}+(34k_{sub}+\log N)N^{\alpha}+6\sqrt{N}\log N}{N-\frac{N^{1-\alpha}}{\epsilon}-5\sqrt{N}\log N}$
$\leq$ $\frac{4N^{1-\alpha}+4(34k_{sub}+\log N)N^{\alpha}+24\sqrt{N}\log N}{N}$
where the second inequality holds because $p_{\mathcal{B}}\leq 1/N$; the third inequality follows from the definition of $N_{2}$, namely $\sum_{n=1}^{N_{2}}\mu_{n}=\lambda N$, which implies $\sum_{n=N_{2}+1}^{N}\mu_{n}=N^{1-\alpha}$. Since $\mu_{n}\geq\epsilon$ for any server $n$, we have $(N-N_{2})\epsilon\leq N^{1-\alpha}$, and hence $N_{2}\geq N-\frac{N^{1-\alpha}}{\epsilon}$; the last inequality holds for a large $N$ such that $\epsilon\geq 2/N^{\alpha}$ and $\sqrt{N}\geq 20\log N.$ Finally, a job that is not routed to an idle server is either blocked or wait in a buffer
$p_{\mathcal{W}}=$ $p_{\mathcal{B}}+p_{\overline{\mathcal{W}}}\leq\frac{4N^{1-\alpha}+4(34k_{sub}+\log N)N^{\alpha}+24\sqrt{N}\log N+1}{N}$
@@
$p_{\mathcal{B}}\leq$ $\mathbb{P}\left(\sum_{n=1}^{N}Q_{n}=Nb\right)\leq 32\left(\frac{2r}{N^{(1-\alpha)}}\right)^{r}.$
@@ Appendix A Tail Bound via Lyapunov Drift Analysis
Remark 1.
Lemma 5 from [32] requires for a nonnegative Lyapunov function but can be more broadly applied to a lower-bounded Lyapunov function. Considering $V(s)$ which is lower bounded and letting $m=\inf_{s\in\mathcal{S}}V(s)$, we can define a shifted function
$W(s)=V(s)-m\geq 0,$
and let D=BmD=B-m. Since $V(s)$ and $W(s)$ differ by a univeral constant for all $s$, the drift remains the same $\nabla W(s)=\nabla V(s),$ and the one-step differences satisfy $|W(s^{\prime})-W(s)|=|V(s^{\prime})-V(s)|$. Therefore, all the conditions of the lemma that hold for $(V,B)$ also hold for $(W,D)$, and the same tail bound follows. Moreover,
$\mathbb{P}\big(W(S(\infty))\geq D+x\big)=\mathbb{P}\big(V(S(\infty))\geq B+x\big),$
so the probability statement is identical. In our analysis, we use lower-bounded Lyapunov functions defined in the most intuitive way without explicitly making them non-negative.
@@ Appendix B Most of the Fastest Servers are Busy
$\mathbb{P}\left(V(q)\geq 1+8\log^{2}N\right)$ $\leq\left(\frac{N}{N+\frac{\delta N}{2}}\right)^{4\log^{2}N}$
@@
$\mathbb{P}\left(V(q)\geq 5\sqrt{N}\log N\right)\leq$ $\left(\frac{N}{N+\mu_{N_{2}}{\sqrt{N}}\log N}\right)^{2\sqrt{N}\log N}$
@@ Appendix C State Space Collapse in Sub-Halfin-Whitt Regime
$=$ $\sum_{n=1}^{N}\lambda N\mathbb{P}_{n}(q)(V(q+e_{n})-V(q))+\mu_{n}\mathbb{I}(q_{n}\geq 1)(V(q-e_{n})-V(q)),$
where $\mathbb{P}_{n}(q)$ denotes the probability that an arriving job joins the server $n$ given the queue state $q.$
@@
    $=$ $\min\left\{\sum_{n=1}^{N}(q_{n}-1)\mathbb{I}(q_{n}\geq 2),~N_{2}+(1-\frac{1}{4\nu})N^{1-\alpha}-\sum_{n=1}^{N}\mathbb{I}(q_{n}\geq 1)-1\right\}-V(q)$
@@
    $\triangledown V(q)\leq$ $-\lambda N+\sum_{m=1}^{N}\mu_{m}\mathbb{I}(q_{m}=1)$
@@
We now apply the tail bound in Lemma 5. As noted in Remark 1, although the Lyapunov function $V(\cdot)$ is not nonnegative, it is lower bounded and can be shifted by a large value (e.g., $N-N_{2}$ in this case) to satisfy the lemma without affecting the drift or the resulting tail bound. Invoking Lemma 5 with $B=c_{0}N^{1-\alpha},$ $\gamma=\frac{\epsilon}{4(b-1)}N^{1-\alpha},$ $\nu_{\max}=1,$ $q_{\max}=N,$ $\delta=N$ and $j=\frac{1-c_{0}}{4}N^{1-\alpha},$ we obtain that for sufficiently large NN
$\mathbb{P}\left(V(Q)\geq\frac{1+c_{0}}{2}N^{1-\alpha}\right)$ $\leq\left(\frac{N}{N+\frac{\epsilon}{4(b-1)}N^{1-\alpha}}\right)^{\frac{1-c_{0}}{4}N^{1-\alpha}}+\left(\frac{N}{\frac{\epsilon}{4(b-1)}N^{1-\alpha}}+1\right)\exp\left(-\epsilon\log^{2}N\right)$